# A First Look at Bugs in LLM Inference Engines

### v1

@@ 7. Discussion
Implications. With the rising popularity of LLM inference engines, our study offers timely and practical insights for different stakeholders.
@@
For inference engine vendors, our taxonomy in Finding 5.1 outlines 28 unique root causes, helping vendors understand and quickly address or avoid these issues. Findings 5.1.15.1.5 highlight areas that require special attention during implementations. For example, Finding 5.1.1 advises developers to concentrate on checking, processing, and error handling for character encoding, tokenizers, and user-specific formatting when dealing with inputs and outputs. While vendors might consider prioritizing different root causes, Finding 4 advises treating all bugs with equal seriousness, since any could pose a high risk of causing crashes. Additionally, developers might wonder which engine components are more prone to errors. This question is answered by Finding 7. For example, for engines targeting cloud environments, special focus should be on the implementation of model loaders and operators.
@@
For LLM app developers, Finding 4.1 highlights the symptoms to consider for fault tolerance during the development of LLM apps, including crashes, unexpected outputs, feature failures, abnormal performance, system hangs, and silent errors. Furthermore, Findings 4.1.14.1.3 recommend stage-specific fault tolerance strategies for LLM app developers. For example, Finding 4.1.1 reveals that 97% of bugs appear as crashes during the setup phase, highlighting the importance of implementing crash fault tolerance during LLM apps setup. Moreover, Finding 6 shows that symptom distribution is consistent across engines, indicating that fault tolerance considerations are the same regardless of the engine used in LLM apps. Additionally, Finding 5.1.4 identifies environmental compatibility as a major (29%) error source, urging LLM app developers to clearly specify target environments and select suitable engines.
For researchers, Finding 4 is the first to empirically identify the features of bugs under LLM undetermined output tokens, including mis-encoded characters, incoherent semantics, inconsistent outputs, abnormal token lengths, and repeated substrings, providing insights for researchers to design oracles. Additionally, Finding 4.2 indicates that researchers should not rely solely on crashes as the oracle for LLM engine testing, as over 35% of bugs exhibit other symptoms. Furthermore, Findings 5.1.15.1.5 offer a foundation for developing more efficient testing methods. For instance, Finding 5.1.2 recommends focusing on configuration combinations, model compatibility, precision, and concurrency to efficiently test bugs under these conditions.
Feedback from the Community. To validate our findings, we have performed a preliminary test on Llama.cpp using our findings. Specially, we employ Finding 5.1.2 to generate configurations composition automatically and Findings 4.2 and 4 to set the ‘crash’, ‘mis-encoded characters’, ‘incoherent semantics’, ‘inconsistent outputs’, ‘abnormal token lengths’, and ‘repeated substrings’ as test oracles. As a result, two bugs [^3] To maintain anonymity, the links of issues here will be added after the paper is accepted. have been confirmed and fixed by the core developers as of this writing. One bug is linked to unexpected output token length under the ‘-n’ configuration, and the other causes a crash under the ‘-ctv’ configuration. This result further demonstrates the validity of our findings.
@@
Threats to Validity. First, our analysis focuses on five LLM inference engines, which may limit the generalizability of our findings to all LLM inference engines. To mitigate this threat, we carefully selected engines with high popularity, well-known providers, and diverse targeted environments. Second, our findings are based on reported issues from GitHub, potentially introducing bias. To mitigate this, we collected a large number of issues, specifically 929 bug-related issues, through careful manual inspection. Third, our methodology involves manual bug categorization, which may introduce bias. To mitigate this, three authors with at least over two years of LLM apps development participate in taxonomy construction, resolving inconsistencies through discussion until reaching an agreement.
@@ 8. Related Work
Empirical Study on Bugs in Machine Learning Libraries. Bugs in machine learning frameworks have long been a significant concern in software engineering (Guo et al., 2019; Chen et al., 2020; Zhang et al., 2018; Chen et al., 2023; Guan et al., 2023; Islam et al., 2019). Chen et al. (Chen et al., 2020) have conducted a comprehensive survey and analysis on Stack Overflow of developers’ questions about deploying machine learning models in the cloud, mobile, and browsers, establishing a taxonomy of specific deployment challenges. Chen et al. (Chen et al., 2023) have conducted an empirical study on four popular deep learning frameworks and measured the test coverage achieved by three testing techniques. While previous studies have investigated bugs in machine learning deployments across various platforms, they primarily focus on traditional machine learning models. Our work focuses on bugs in LLM inference engines, which have unique characteristics, such as undetermined output, high resource demands, and compatibility for diverse environments.
Testing of Machine Learning Libraries. Many studies have proposed different methods to test machine learning frameworks, including compilers (Liu et al., 2023, 2022; Ma et al., 2023; Luo et al., 2021) and libraries (Gu et al., 2022; Luo et al., 2021; Wang et al., 2020; Li et al., 2023; Pham et al., 2019; Yang et al., 2023; Wei et al., 2022; Deng et al., 2023, 2024). Liu et al. (Liu et al., 2023) have proposed NNSmith, which leveraged symbolic constraint solving and gradient-based search to generate diverse yet valid DNN test models for fuzzing deep learning compilers. Additionally, some studies leverage large language models for testing deep learning libraries (Deng et al., 2023, 2024). TitanFuzz (Deng et al., 2023) leverages Large Language Models (LLMs) for generating valid input programs to fuzz deep learning libraries. Although existing work has explored various components of DL frameworks, no studies have specifically addressed the bugs in LLM inference engines. Our work provides an empirical foundation for future research on LLM inference engines. Moreover, our open-source dataset can serve as a data source for enhancing the LLM-based testing methods.
@@ 9. Conclusion
We present the first study on bugs in LLM inference engines, addressing a critical gap in understanding the symptoms, root cause, and commonality of bugs. By examining five well-known engines, we identify six bug symptoms, 28 root causes, and strong correlation of these characteristics across engines, offering insights into efficient bug detection and location. Our primary findings include that (1) Five factors are useful for oracle setting under undetermined LLM outputs; (2) Diverse oracles beyond crashes are needed for LLM inference engine testing; (3) We reveal actionable diagnostic patterns for all identified bug symptoms. Our research provides an empirical foundation for analyzing and testing LLM inference engines, with practical implications for different stakeholders.

### v2

@@ 10.2. Implications
With the rising popularity of LLM inference engines, our study offers timely and practical insights for different stakeholders.
@@
For inference engine vendors, our taxonomy in Finding 5.1 outlines 28 unique root causes, helping vendors understand and quickly address or avoid these issues. Findings 5.1.15.1.5 highlight areas that require special attention during implementations. For example, Finding 5.1.1 advises developers to concentrate on checking, processing, and error handling for character encoding, tokenizers, and user-specific formatting when dealing with inputs and outputs. While vendors might consider prioritizing different root causes, Finding 5.3 advises treating all bugs with equal seriousness, since any could pose a high risk of causing crashes. Additionally, developers might wonder which engine components are more prone to errors. This question is answered by Finding 6. For example, for engines targeting cloud environments, special focus should be on the implementation of model loaders and operators.
@@
For LLM app developers, Finding 4.1 highlights the symptoms to consider for fault tolerance during the development of LLM apps, including crashes, unexpected outputs, feature failures, abnormal performance, system hangs, and silent errors. Furthermore, Findings 4.2.14.2.3 recommend stage-specific fault tolerance strategies for LLM app developers. For example, Finding 4.2.1 reveals that 97% of bugs appear as crashes during the setup phase, highlighting the importance of implementing crash fault tolerance during LLM apps setup. Moreover, Finding 6 shows that symptom distribution is consistent across engines, indicating that fault tolerance considerations are the same regardless of the engine used in LLM apps. Additionally, Finding 5.1.4 identifies environmental compatibility as a major (29%) error source, urging LLM app developers to clearly specify target environments and select suitable engines.
For debugging memory-related bugs, our study provides a key diagnostic heuristic: Memory problems (e.g., memory leaks, OOM, abnormal memory growth) may not solely originate from bugs in resource management modules. Rather, due to the immense resource demands of LLM inference, incorrect functional logic can lead to excessive resource allocation during implementation or a failure to deallocate resources after an operation is complete. For developers, this suggests a strategy of investigating both the logic layer and the resource management layer in parallel when encountering memory-related bugs. While inspecting the memory allocator or KV cache scheduler, developers should place a strong emphasis on investigating the functional implementation, such as custom CUDA kernels, complex attention mechanisms, or sampling algorithms. For instance, a bug is more likely to be an incorrect tensor shape calculation (RC.3 Mismatched Shape) causing an out-of-bounds write, or a flawed algorithm (RC.1 Incorrect Algorithm Implementation) that fails to release memory on an edge case, than a bug in the memory pool manager itself.
For researchers, our study provides a foundational empirical basis that opens several specific and actionable research directions. Finding 4.2.3 is the first to empirically identify the features of bugs under LLM undetermined output tokens, providing concrete insights for researchers to design test oracles that move beyond simple crash detection. For instance, these factors can be translated into concrete automated oracles. Inconsistent outputs can be detected via a reproducibility oracle that runs the same prompt multiple times under deterministic decoding (e.g., greedy decoding with fixed seeds) and asserts identical outputs. Abnormal length and repetition can be flagged by simple checks, such as empty or single-token outputs, max-token exhaustion, high repeated $n$-gram rates (e.g., n=3n{=}355), or consecutive substring repetition. Mis-encoded characters can be identified by validating that outputs are well-formed Unicode/UTF-8 and by searching for the Unicode replacement character (e.g., U+FFFD) or unexpected control characters. For judging incoherent semantics, more advanced methods can be explored, such as employing a powerful external LLM as a judge to evaluate the semantic coherence of the generated text. These examples demonstrate how our findings enable the development of practical, automated techniques to validate LLM outputs at scale. Additionally, Finding 4.2.3 indicates that researchers should not rely solely on crashes as the oracle for LLM engine testing, as over 35% of bugs exhibit other symptoms, underscoring the need for a diverse set of oracles that encompass performance anomalies, feature failures, and silent errors to achieve comprehensive testing coverage.
Furthermore, Findings 5.1.15.1.5 offer a robust foundation for developing more efficient and targeted testing methods by highlighting key vulnerability areas. For instance, Finding 5.1.2 reveals that many bugs are triggered under complex configuration conditions, such as model compatibility, mixed precision, and concurrency, highlighting the critical need for configuration-aware testing. Researchers could design advanced configuration-space exploration techniques specifically for LLM inference engines, such as combinatorial interaction testing or adaptive sampling, to efficiently navigate the high-dimensional parameter space of LLM engines, thereby improving test efficiency by focusing on high-risk combinations that are prone to errors. Similarly, Findings 5.1.3 and 5.1.4 point to the critical need for testing environmental compatibility and resource management, advocating for techniques like static analysis for resource leak detection and metamorphic testing across backends. By addressing these engine-specific challenges, researchers can advance the state of testing for LLM inference systems, ensuring their reliability and performance across diverse deployment scenarios.
Moreover, the strong commonality of bug patterns across different engines, as evidenced by Finding 6, indicates a significant level of cross-engine generalizability. This suggests that researchers may investigate whether oracle definitions and testing techniques, such as those based on output validation or configuration testing, can be generalized across inference engines, enabling the creation of reusable research artifacts. For example, standardized test suites or metamorphic relations derived from one engine could be adapted to others, reducing duplication of effort and accelerating the validation of new or updated engines. By leveraging these shared patterns, researchers can develop more scalable and portable methods for ensuring the reliability of LLM inference systems in diverse environments.
10.3. General Guidelines for Developing LLM Inference Engines
Based on our findings, we aim to provide general guidelines for developing LLM inference engines structured by the composited modules.
Engine Setup Module. The engine setup phase is critical for ensuring a stable foundation, as bugs at this stage often prevent deployment entirely. (1) LLM engine developers can conduct thorough cross-platform testing across all supported backends, platforms, and devices, as Finding 5.1.4 implies that 29% of issues arise from these factors. For instance, continuous integration pipelines should include targeted testing on heterogeneous CUDA environments (vLLM issue#7472) and the Metal backend of Apple Silicon (Llama.cpp issue #1737), validating core functions like tensor operations and memory management. (2) Automated dependency checkers can be employed to ensure compatibility of dependencies and installation configurations, as Finding 5.1.2 indicates that incorrect setup configuration is a major root cause. For example, these checkers should be designed to validate critical configuration aspects, such as ensuring consistency in parallelism configurations across distributed settings (DeepSpeed issue #3628), maintaining precision alignment between models, engines, and inputs (TensorRT-LLM issue #1303), verifying the completeness of dependencies for specific deployment modes (vLLM issue #8107), and checking the availability of system libraries and headers during source compilation (Llama.cpp issue #279). (3) Developers can prioritize detecting crash symptoms because Finding 4.2.1 suggests most bugs in this phase lead to crashes.
Model Conversion Module. (1) LLM engine developers can implement validation steps during and after conversion (e.g., checksums and precision checks) to ensure converted models are complete and configured correctly, as Finding 4.2.2 suggests that generated model files can be flawed. For example, validators should check for complete parameter sets and correct precision after AWQ quantization (TensorRT-LLM issue #1245) to avoid silent errors during inference. Moreover, developers can establish a comprehensive testing matrix covering critical configuration combinations. This matrix should systematically evaluate interactions between model architectures and formats(e.g., TensorRT-LLM issue #1634), quantization methods and precision settings (e.g., testing INT4, AWQ, and GPTQ across different precision backends), and target backend platforms (e.g., validating CUDA, Metal, and Vulkan conversions for each model type). Such combinatorial testing can proactively identify incompatibilities that lead to crashes, incomplete outputs, or misconfigured models. (2) Given that Finding 6 indicates model conversion tools are highly bug-prone, developers can regularly update and test these tools for their compatibility, correctness, and performance. For instance, test suites should cover edge cases like tensor shape mismatches during SmoothQuant conversion (TensorRT-LLM issue #899) and validate output model compatibility across backends.
Inference/Serving Module. (1) LLM engine developers can implement multiple testing oracles, as suggested by Findings 4.2.3 and 4.2.3. For example, validation of output tokens can be incorporated to check features described in Finding 4.2.3 (e.g., character correctness, semantic coherence, repetition). (2) Because Finding 4.1 implies that abnormal performance is a major bug symptom in this phase, developers can conduct performance tests under different resource settings. For example, performance tests should measure throughput and latency under heterogeneous LLMs (DeepSpeed issue #2488) and validate KV cache management under memory constraints. (3) Building on Finding 5.1.2, which reveals that configuration-related bugs comprise 23% of all issues, it is essential to formulate a rigorous testing strategy that specifically targets high-risk configuration combinations in the inference/serving stages. This strategy should center on designing comprehensive testing matrices to evaluate critical interactions across multiple dimensions, including concurrency settings, resource limits, job parameters, model types, backend devices, and parallel configurations. For example, a focused approach involves testing various worker counts under constrained KV cache sizes to preempt shared state conflicts with concurrent workers (Llama.cpp issue #3537). This should be extended to experimenting with diverse job parameters, such as max_tokens values, sampling options, and grammar constraints applied across a range of model architectures. Additionally, validation must ensure compatibility in multi-GPU environments with tensor parallelism enabled (vLLM issue #7472), and verify the correct synchronization of parallel workers to prevent performance anomalies.
10.4. Threats to Validity
@@
First, our analysis focuses on five LLM inference engines, which may limit the generalizability of our findings to all LLM inference engines. To mitigate this threat, we carefully selected engines with high popularity, well-known providers, and diverse targeted environments. Second, our findings are based on reported issues from GitHub, potentially introducing bias. To mitigate this, we collected a large number of issues, specifically 929 bug-related issues, through careful manual inspection. Third, our methodology involves manual bug categorization and the two inspectors involve in the coding process are doctoral students, which may introduce bias. To mitigate this, three authors with at least over two years of LLM apps development participate in taxonomy construction, resolving inconsistencies through discussion until reaching an agreement. Moreover, the final Kappa values for the symptom and root cause labeling are greater than 0.8, indicating almost perfect agreement and confirming the reliability of the coding schema. Fourth, this study focuses only on closed issues to ensure sufficient evidence for reliable characteristic labeling and analysis. The trade-off of this methodological choice is that we may omit newly reported or unresolved problems in the open issues, which might include recent challenges as LLM inference engines evolve rapidly. Future work could complement this study by incorporating open issues once they are resolved or better documented.
@@ 11. Related Work
Empirical study on bugs in machine learning libraries. Bugs in machine learning systems have long been a significant concern in software engineering (Guo et al., 2019; Chen et al., 2020; Zhang et al., 2018; Chen et al., 2023; Guan et al., 2023; Islam et al., 2019). Chen et al. (Chen et al., 2020) have conducted a comprehensive survey and analysis on Stack Overflow of developers’ questions about deploying machine learning models in the cloud, mobile, and browsers, establishing a taxonomy of specific deployment challenges. Chen et al. (Chen et al., 2023) have conducted an empirical study on four popular deep learning frameworks and measured the test coverage achieved by three testing techniques. While previous studies have investigated bugs in machine learning deployments across various platforms, they primarily focus on traditional machine learning models. Our work focuses on bugs in LLM inference engines, which have unique characteristics, such as undetermined output, high resource demands, and compatibility for diverse environments.
Testing of machine learning libraries. Many studies have proposed different methods to test machine learning frameworks, including compilers (Liu et al., 2023b, 2022; Ma et al., 2023; Luo et al., 2021) and libraries (Gu et al., 2022; Luo et al., 2021; Wang et al., 2020; Li et al., 2023; Pham et al., 2019; Yang et al., 2023; Wei et al., 2022; Deng et al., 2023, 2024). Liu et al. (Liu et al., 2023b) have proposed NNSmith, which leveraged symbolic constraint solving and gradient-based search to generate diverse yet valid DNN test models for fuzzing deep learning compilers. Additionally, some studies leverage large language models for testing deep learning libraries (Deng et al., 2023, 2024). TitanFuzz (Deng et al., 2023) leverages Large Language Models (LLMs) for generating valid input programs to fuzz deep learning libraries. Although existing work has explored various components of DL frameworks, no studies have specifically addressed the bugs in LLM inference engines. Our work provides an empirical foundation for future research on LLM inference engines. Moreover, our open-source dataset can serve as a data source for enhancing the LLM-based testing methods.
@@ 12. Conclusion
We present the first study on bugs in LLM inference engines, addressing a critical gap in understanding the bug symptoms, root cause, commonality, fix effort, fix strategies, and temporal evolution of bugs. By examining five well-known engines, we identify six bug symptoms, 28 root causes, and strong correlation of these characteristics across engines, offering insights into efficient bug detection and location. Our primary findings include that (1) Five factors are useful for oracle setting under undetermined LLM outputs; (2) Diverse oracles beyond crashes are needed for LLM inference engine testing; (3) We reveal actionable diagnostic patterns for all identified bug symptoms. Our research provides an empirical foundation for analyzing and testing LLM inference engines, with practical implications for different stakeholders and general guidelines for LLM inference engine development.
Acknowledgment
This work was supported by Beijing Natural Science Foundation (No. L253005), the National Natural Science Foundation of China (No. 62325201), the Key Laboratory of High Confidence Software Technologies (Peking University), the Ministry of Education, and the Center for Data Space Technology and System, Peking University.